OpenStudy (anonymous):

Solve this equation... ∫xe^-x^2 dx??? I WILL GIVE OUT A MEDAL TO WHOEVER CAN HELP!

OpenStudy (zzr0ck3r):

u=-x^2 du=-2x dx

OpenStudy (anonymous):

Only one person gets a medal sorry it won't let me do more than one.

OpenStudy (anonymous):

Troll????

OpenStudy (zzr0ck3r):

so \(\frac{-1}{2}\int e^u du=-\frac{1}{2}e^u+c=-\frac{1}{2}e^{-x^2}+c\)

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

Why are you calling me a troll you idiot?

OpenStudy (zzr0ck3r):

dont worry about them, focus on me:) do you understand what I did?

OpenStudy (anonymous):

Yes. kinda.

OpenStudy (anonymous):

You called me a troll

OpenStudy (anonymous):

No i did not..

OpenStudy (anonymous):

on my page you did

OpenStudy (anonymous):

DID NOT. I DONT EVEN KNOW YOU. WTF..

OpenStudy (zzr0ck3r):

let \(u=-x^2\) then \(\frac{du}{dx}=-2x\) so \(dx=\frac{du}{-2x}\) plug this into your integral \(\int xe^{-x^2}dx=\int xe^u \frac{du}{-2x}=-\frac{1}{2}\int e^udu=-\frac{1}{2}e^u+c\) plugging back in for \(u\) we get \(\int xe^{-x^2}dx=-\frac{1}{2}e^{-x^2+c}\)

OpenStudy (zzr0ck3r):

that should say \(\int xe^{-x^2}dx=-\frac{1}{2}e^{-x^2}+c\)

OpenStudy (zzr0ck3r):

do you understand?

OpenStudy (anonymous):

Yup :)